### v1

# Simulating Classroom Education with LLM-Empowered Agents

Abstract
Large language models (LLMs) have been employed in various intelligent educational tasks to assist teaching. While preliminary explorations have focused on independent LLM-empowered agents for specific educational tasks, the potential for LLMs within a multi-agent collaborative framework to simulate a classroom with real user participation remains unexplored. In this work, we propose SimClass, a multi-agent classroom simulation framework involving user participation. We recognize representative class roles and introduce a novel class control mechanism for automatic classroom teaching, and conduct user experiments in two real-world courses. Utilizing the Flanders Interactive Analysis System and Community of Inquiry theoretical frame works from educational analysis, we demonstrate that LLMs can simulate traditional classroom interaction patterns effectively while enhancing user’s experience. We also observe emergent group behaviors among agents in SimClass, where agents collaborate to create enlivening interactions in classrooms to improve user learning process. We hope this work pioneers the application of LLM-empowered multi-agent systems in virtual classroom teaching.
Simulating Classroom Education with LLM-Empowered Agents
Zheyuan Zhang♠††thanks: Equal contribution, Daniel Zhang-Li♠11footnotemark: 1, Jifan Yu♡, Linlu Gong♠, Jinchang Zhou♠, Zhiyuan Liu♠, Lei Hou♠, Juanzi Li♠ ♠Department of Computer Science and Technology; ♡Institute of Education Tsinghua University, Beijing, 100084, China zheyuan-22@mails.tsinghua.edu.cn lijuanzi@tsinghua.edu.cn
1 Introduction
The pursuit of utilizing artificial intelligence to provide immediate and customized teaching for students origins from the era of Intelligent Tutoring Systems (ITS) Nwana (1990). Following this enthusiasm, from personalized educational recommendation systems Liu et al. (2019) to teaching assistants Tu et al. (2023); Khan Academy (2024) and even LLM-driven AI teacher Markel et al. (2023); Yue et al. (2024), researchers have conducted enormous technological explorations and achieved impressive performance in specific educational tasks.
As technology advances, intense discussions have also emerged around this topic concerning methodologies Extance (2023); Yue et al. (2024). One of the most central directions is how to fully leverage the capabilities of large models to simulate real classrooms with multiple agents for automated teaching. From an educational perspective, this approach allows large models to move beyond their instrumental use and delve deeper into educational paradigms Lave (1996); Opara et al. (2023). From a technical standpoint, multi-agent collaboration technologies Qian et al. (2024) could further stimulate the latent knowledge of large models in education, leading to the emergence of richer capabilities Li et al. (2024a); Aher et al. (2023).
However, towards LLM-empowered multi-agent systems that involve real user participation, there are still several fundamental research questions that need to be explored. (1) Simulation Capability Assessment: To what extent can a multi-agent classroom powered by large models simulate real teacher-student interactions? (2) Learning Experience Measurement: Can students in such an intelligent teaching environment experience a high sense of presence and learn effectively? (3) Emergence Phenomenon Observation: What types of classroom behaviors may spontaneously arise in scenarios that integrate multiple agents?
In this work, responding to the questions above, we present SimClass, a Multi-Agent Classroom Simulation framework, and conduct real-world observation along with analysis based on it. To better simulate the classroom, we recognize representative class roles and design a novel class control mechanism with functional workflows. For systematic experiments, we deploy 2 different courses with prepared slides and teaching scripts as basis. 48 students are invited to join the classroom, learning and interacting with the system, and all the behavioral data is carefully recorded. Then we conduct experiments to explore the mentioned questions. (1) Firstly, we apply the Flanders Interaction Analysis System Amatari (2015) to evaluate the interactions happening in the SimClass and explore the interaction pattern of the agents’ classroom. (2) Secondly, we analyze the educational experience of these users, particularly with Community of Inquiry theory Garrison and Arbaugh (2007). (3) Finally, we summarize several emergent group actions during the experiment for qualitative analysis.
During our experiments, we observe the effectiveness of the class role and control mechanism design. Based on the problems we identified, experimental results show that: (1) Similarity: SimClass exhibits behaviors, interaction patterns, and characteristics similar to those of traditional classrooms; (2) Effectiveness: Multiple classroom agents enable users to engage more effectively in class and enhance their sense of presence; (3) Emergence: Our control mechanism spontaneously elicits the emergent behaviors in the multi-agent classroom system, including collaborative teaching and discussion, emotional company and discipline control. In summary, the LLM-based multi-agent system demonstrates the potential for simulating real classroom environments for educational purposes. We hope our work serves as a pioneering effort in this direction. The dataset of classroom interactions between users and multiple LLMs will be released soon for both education and AI researchers.
2 Related Work
2.1 LLMs for Human Simulation
Recently, Large Language Models (LLMs) have achieved remarkable breakthroughs in various natural language processing (NLP) tasks Brown et al. (2020); OpenAI (2024); Touvron et al. (2023); Team (2024). The intelligence they demonstrated opened up opportunities and possibilities for applications in many other scenarios Bubeck et al. (2023); Yang et al. (2023). As LLMs encode many human-like behaviors in their training data, an increasing number of researchers are utilizing LLMs for human scenario simulation, investigating the model’s capabilities for decision and actions as LLM-Empowered Agents in many fields, such as social and psychological research Aher et al. (2023); Park et al. (2023); Li et al. (2024a); Gao et al. (2023); Li et al. (2024d); Zhang et al. (2024), software development Qian et al. (2024); Hong et al. (2023), chemical and medicine Li et al. (2024c); M. Bran et al. (2024), and games Wang et al. (2023). Novel collaboration techniques are explored to enhance the cooperation and performance of multi-agent systems Cheng et al. (2024); Wu et al. (2023). These works offer technical possibilities for multi-agent education and inspire curiosity about potential emergent phenomena.
2.2 LLMs for Education
With the eminent linguistic capabilities, explanatory skills, and parameterized knowledge of LLMs, numerous studies have explored applying LLMs to education services. In addition to applying large models to downstream tasks in the education Hu et al. (2024); Li et al. (2024b); Jeon and Lee (2023), many researchers are applying these models to replace certain classroom aspects, such as playing students to train teachers Lee et al. (2023); Markel et al. (2023) or playing instructors to teach students Tu et al. (2023); Sonkar et al. (2023); Khan Academy (2024); Chen et al. (2023). Yue et al. (2024) explored the use of multiple student agents to assist students in discussion, though they haven’t involved real users. Existing work has examined various facets of interactions between LLMs and humans in educational settings.
3 SimClass
3.1 Overview
The design principles for constructing this immersive simulated classroom originate from the following two concerns: (1) How to ensure that the classroom covers the core teaching behaviors? (2) How to maintain the entirety of the interaction within the natural flow of the classroom process?
For the former concern, we categorize classroom interaction behaviors based on widely accepted pedagogy principles Schwanke (1981): Teaching and Initiation (TI), the teacher’s teaching and the feedback or ideas expressed by students; In-depth Discussion (ID), alignment, discussion, and multiple Q&A between teacher and students to help students construct understanding of concepts; Emotional Companionship (EC), encouraging students to learn, creating a positive learning atmosphere, and providing emotional support; and Classroom Management (CM), maintaining discipline, organizing disruptive behaviors, and guiding the classroom content. Given that these behaviors are realized through the varied Class Roles (denoted as , where each denotes a certain role), it is essential to ensure the diversity and coverage of proposed agents within the classroom.
For the latter concern, we need to ensure that the interactions among multiple agents within the system are finely and rhythmically controlled within the course content. Given the Learning Materials (denoted as , where each teaching script is organized by order), we propose a novel Session Controller to manage the course interaction flow based on class status and the help of a core manager agent Wu et al. (2023).
Based on these principles, we construct multiple class roles, implement class control, and ultimately derive the simulated classroom process.
3.2 Class Role Agentization
The teaching and learning process is presented as an informative, multi-round, and task-oriented communication Lave (1996). However, simply exchanging responses of LLMs inevitably faces significant challenges including role flipping, instruction repeating, and fake replies Qian et al. (2024). Consequently, following the classroom behaviors outlined previously, we define two types of agents: Teaching Agents and Classmate Agents. Each agent is facilitated through prompting LLMs and associated with one or more class roles, denoted as:
where is the role customization operation, is the system prompt with agent description.
Teaching Agents
The teacher and the teaching assistant are the authoritative party responsible for imparting knowledge in the classroom, encompassing most teaching behaviors. The acronyms in parentheses represent the roles that the agent needs to accomplish in a classroom environment.
Teacher Agent (TI, ID, EC, CM) : Given the teaching scripts , its task is to persuasively display material to students or answer questions based on the classroom historical discussions .
Assistant Agent (ID, EC, CM): Given the classroom history , the assistant is responsible to supplement teaching information, participate in discussion, maintain the discipline and continuity of the class, and enhance student learning efficiency.
Classmate Agents
This type of agents are incorporated in addition to the teaching agents with distinct personality traits to better simulate traditional one-to-many classrooms, performing peer student roles. In this paper, we initialize typical classmates, while users can also freely customize and deploy more interesting classmate agents on the platform.
Class Clown (TI, EC, CM): This agent is designed to initiate ideas, enliven the atmosphere, help the user as a peer, and help the teachers to guide the class direction when the user is distracted.
Deep Thinker (TI, ID): This agent aims to do deep thinking and raise topics that challenge the knowledge of the classroom.
Note Taker (TI, CM): This agent loves to summarize and share notes for classroom content, helping everyone to organize their thoughts.
Inquisitive Mind (TI, EC): This agent frequently poses questions about lectures, which stimulates others’ thinking and discussion.
3.3 Classroom Session Controller
Unlike Standardized Operating Procedures (SOPs) multi-agent systems Qian et al. (2024); Hong et al. (2023), the classroom scenario is a dynamic group chat without a strict workflow, where agents need to dynamically determine the appropriate speaking timing. Therefore, we implement a controller that observes, makes decisions, and controls agents to behave based on the current Class State. The Session Controller includes the following modules: Class State Receptor, Function Executor, and Manager Agent.
Class State Receptor
Let the classroom dialogue history until time denote as , where is the utterance posted by agent or user (denoted as ). The class state is composed as:
where is composed of the learning materials that have been taught until .
Functions
We design and divide the actions in the classroom into a functional hierarchy with two major categories. Tutoring functions can only be performed by teacher agent , such as teaching by displaying scripts and going to the next material page . Interacting functions can be performed by each agent . According to the context, the interaction will emerge as diverse classroom activities, which are discussed in subsequent experiments. These functions are pluggable, allowing the addition of newly defined functions for different agents, such as displaying exercises.
Manager Agent
Following AutoGen Wu et al. (2023) and MathVC Yue et al. (2024), we design a hidden and meta agent to regulate the speakers. This agent receives the current class state , observes and understands the class process, and decides the next action to be executed.The task of Manager Agent can be defined as:
where is a certain kind of function, and the action will be executed and refresh the whole class into the next state. Specifically, the system will wait for a time window after an action is performed. If the user speaks or the waiting period ends, it will trigger the manager agent to make a new decision.
3.4 Classroom Demonstration
After introducing the necessary component of the SimClass , we present the demonstration of an entire class process: (1) Initialization. At the beginning of the class, the first function will be executed, displaying the initial course script and slides. At this point, users can interact with the class, and the manager agent will start controlling the class flow; (2) Tutoring and Interaction: the manager agent will continuously observe and control the class based on the states, and other agents will perform diverse activities by collaboration. As the example shown in Figure 1, when a user asks about the course content, the classroom interaction flow may involve the assistant responding, the teacher providing additional information, and sometimes the classmate agents raising corresponding topics; (3) Ending. After all the learning materials are taught and the final discussion ends, the classroom will close and provide survey questions to users.
4 Experiments
To evaluate the performance of SimClass, we invite a group of university students to participate in the classroom to record interaction data and collect feedback from them. We also develop ablation systems to better understand the impact of various interaction types within SimClass. Our analyses mainly focus on three key aspects: classroom interactions, user experience, and the emergent group behaviors of the LLM-empowered agents.
4.1 Experimental Setup
Courses and Materials. We conduct experiments with two courses. Two experienced teachers are invited to design the slides and teaching scripts for the courses. The first one, TAGI, Towards Artificial General Intelligence, covers the development of AI and language models, consisting of 50 pages of slides, each with a corresponding teaching script. The second course, HSU, How to Study at University, addresses topics such as completing academic work, managing pressure, communicating with others, and achieving self-fulfillment, and includes 45 pages of slides and teaching scripts.
Systems. We use GPT-4 as the backbone LLM of both Class Roles and Manager Agent in SimClass. Besides, we implement two ablation systems to investigate the effects of different interaction types in the classroom. In the first system (SimClass w/o classmates), the classmate agents are removed, and only teacher agents are present. In the second system (SimClass w/o interactions), both classmate agents and user input are disabled, resulting in no interaction at all. The teacher can only conduct lectures persistently, and the Manager Agent is limited to the tutoring function.
Participants. In the experiment, we invite 48 university students from different majors to participate in learning two courses on SimClass, with each course and each setting involving 8 users. To ensure the quality control of course data, we invite course designers to create four questions for each course and require users to complete these questions after finishing the course. Data from participants with an accuracy of 50% or lower on these questions were excluded. Ultimately, data from 38 participants remained, with each course and setting having data from at least 5 users. Participants are informed that all course data is generated by AI and needed to be carefully discerned. Each participant receives the appropriate amount of compensation.
| Please rate the overall performance of the platform: |
| Cognitive Presence |
| Does the platform help students to understand concepts and master the corresponding knowledge? |
| Teaching Presence |
| Does the class as a whole serve a specific instructional goal, aligning with the course design and direction? |
| Social Presence |
| Can the responses create a credible and engaging interactive environment in the classroom, encouraging students to participate in interactive learning? |
Survey Content. In addition to the four test questions, each participant is required to complete a short survey composed of three questions regarding the experience after finishing the course. We apply the widely recognized Community of Inquiry (CoI) theory Garrison et al. (1999) in online learning to evaluate the experience of students. Specifically, we adapt the three key elements from CoI to measure the learning experience on SimClass: Cognitive Presence, the degree to which learners are able to construct and confirm meaning through sustained reflection and interaction; Teaching Presence, the extent to which the class is focused, designed, and planned with specific directions and learning objectives; and Social Presence, the ability of learners to project themselves socially and emotionally within a group Garrison and Arbaugh (2007). Students are asked to rate the system on a scale of [0,1,2], with a higher score indicating better performance according to detailed guidelines. The survey questions are listed in Table 1, and further details can be found in Appendix A.
4.2 Statistical Results
| Course | Teacher | Assistant | Students | User |
| TAGI | 353.0 | 82.3 | 123.0 | 18.9 |
| - w/o stu. | 358.2 | 71.1 | - | 13.9 |
| - w/o int. | 398.8 | - | - | - |
| HSU | 218.3 | 90.6 | 147.7 | 15.5 |
| - w/o stu. | 212.3 | 68.2 | - | 8.2 |
| - w/o int. | 228.5 | - | - | - |
Table 2 presents the average speech length of various roles and users across different settings. All systems employ the same teaching scripts, leading to the teacher’s speech being the longest and most closely aligned with the scripts. The assistant’s primary role is to maintain discipline, resulting in shorter dialogues. Classmate agents are generally more talkative, whereas users tend to use fewer words. Notably, the absence of classmate agents significantly reduces the speech length of users and assistant in both courses. The presence of classmate agents in the classroom appears to encourage users to engage in longer conversations.
4.3 Interaction Analysis
To understand the dynamics of SimClass as a multi-agent classroom system, we encode classroom activities into quantitative behaviors. We utilize the Flanders Interaction Analysis System (FIAS) Amatari (2015), a valuable tool for analyzing the verbal behaviors in traditional classrooms. We adapt the method to our simulated classroom system, SimClass, where interactions occur in natural language.
| Speaker | Type | Action |
| Teacher | Indirect Influence (Response) | 1. Accept Feelings |
| 2. Praises or Encourages | ||
| 3. Accept Ideas | ||
| 4. Ask Questions | ||
| Direct Influence (Initiation) | 5. Lecturing | |
| 6. Giving Direction | ||
| 7. Criticizing | ||
| Student | Response | 8. Response |
| Initiation | 9. Initiation | |
| Silence | Silence | 10. Silence or Confusion |
Encoding the Interactions. As shown in Table 3, the FIAS categorizes interactions into ten distinct types: seven for teachers, two for students, and one for silent. Labels represent Indirect Influence from the teacher, while labels indicate Direct Influence. When classroom activities are encoded as sequences, the proportion of each interaction type and their transitions can be decoded to reveal the classroom style, teaching style, and other features. For the classroom history of each student, we prompt GPT-4 to label interactions according to the ten communication categories. We assess the quality of GPT-4’s labeling in Appendix B. The classroom interactions are encoded as sequences, and the two-step transitions of classroom activities are recorded in a matrix . Following the method introduced by Amatari (2015), we add a (silence) to the beginning and end of each class sequence and sum the matrices of students in the same setting to provide a general view of the interactions: . To interpret the classroom interaction Matrix and observe features in SimClass, we report the following metrics designed by Amatari (2015):
Teacher Talk (TT) and Student Talk (ST). TT and ST represent the proportions of total tallies in specific categories that indicate the amount of talk from teacher and students. Respectively, TT and ST are calculated using categories and .
ID Ratio (IDR). This ratio measures the balance between a teacher’s indirect and direct methods of communication and teaching in the classroom. It is calculated by dividing the sum of tallies in categories (Indirect influence) by the sum of tallies in categories (Direct influence).
Student Initiation Ratio (SIR). SIR evaluates the extent to which students initiate interactions themselves during classroom activities, which measures how much students are actively engaging in the classroom. It is calculated by dividing the tallies in category by the total tallies in categories .
| Course | TT | ST | IDR | SIR |
| TAGI | 0.824 | 0.162 | 0.130 | 0.493 |
| - w/o stu. | 0.864 | 0.121 | 0.119 | 0.446 |
| - w/o int. | 0.980 | 0.000 | 0.005 | 0.000 |
| HSU | 0.847 | 0.136 | 0.127 | 0.597 |
| - w/o stu. | 0.896 | 0.087 | 0.164 | 0.233 |
| - w/o int. | 0.978 | 0.000 | 0.089 | 0.000 |
Results. Figure 2 presents the FIAS matrices of SimClass for TAGI and HSU courses. Each matrix is divided into four parts based on the type of interaction in the class, labeled as follows: A (top left): Interactions from teacher to teacher; B (top right): student or silence to teacher; C (bottom left): teacher to student or silence; and D (bottom right): student to student. The matrices reveal the following findings: (1) In part A, most teacher actions are associated with lecturing (Cat. ), where teachers primarily give lectures and interact with the class; (2) Part B demonstrates the teacher’s responses to students. When students initiate ideas or responses to teachers, the teacher praises (Cat.), accepts their ideas (Cat.), or continues teaching; (3) Part C shows student actions in response to teachers, where students mostly initiate questions or respond to lectures; (4) Part D shows that student-to-student interactions and discussions occur periodically. The results of the ablation systems, demonstrated in Appendix C, indicate that interactions are much less diverse after ablation.
Table 4 presents the metric results of FIAS. TT and ST represent the proportion of teacher and student speaking time, respectively. In traditional classrooms, the statistics for TT and ST range from 77.1%-82.6% and 17.4%-22.9% (excluding silence) Zhang et al. (2023). SimClass exhibits a similar distribution. The IDR is low, which is partly due to the higher proportion of script-based teaching. The SIR is relatively high, especially in scenarios involving classmate agents, where there are more instances of students initiating questions.
Conclusion of Interactive Analysis. From the perspective of ratio analysis (Table 4) and interaction distribution (Figure 2), SimClass demonstrates the characteristics of the traditional classroom, effectively simulates traditional classrooms and has the potential to achieve the performance of real classrooms. We further investigate the user experience and illustrate a few in the following sections.
4.4 User Experience
In this section, we report the results from the student experience with SimClass. As shown in Figure 3, several key findings are observed: (1) Importance of Interactions. Interactions during class are crucial for users. Without interaction, user experience significantly declines across all three metrics. (2) Enhancement by Classmate Agents. Classmate agents enhance user experience in terms of Cognitive Presence and Social Presence. This enhancement may be attributed to the classmate agents’ active engagement in asking questions to the teacher, which aids the user’s understanding of concepts and increases the sense of Social Presence in the classroom. (3) Satisfying Teaching Presence. All systems demonstrate good Teaching Presence, maintaining a focused and coherent class. This metric largely depends on the quality of the teaching scripts used, though we observe that interaction and Student Roles slightly improve the user experience. (4) Better Experience in HSU with students. The HSU course achieves a better user experience with the full setting. HSU focuses on college interpersonal relationships and learning methods, where peer learning plays a more significant role. This suggests that a multi-agent design is particularly crucial for certain types of classes.
Conclusion of user experiments. According to Figure 2 and Figure 3, the classroom demonstrates the effectiveness in SimClass in terms of both interactions and user presence.
| Teaching and Initiation |
| [Teacher]: Let’s talk about Alan Turing. … The concept of a thinking machine he proposed essentially foresaw many characteristics of modern intelligent systems, including the ability to learn, adapt, and handle complex problems. |
| [Classmate]: I wonder how Turing would react if he were alive today and saw the voice assistants on our smartphones. But he might also remind us that true intelligence is not merely about mimicking human behavior; it must involve its own “thinking” and “understanding” … |
| In-depth Discussion |
| [Teacher]: … other challenges include the “vanishing and exploding gradients problem” … |
| [User]: What is that? Will it really explode? |
| [Teacher]: It’s not actually explode, but I’m glad you’re interested in the concept! “Vanishing” and “exploding” refer to numerical issues during backpropagation. … |
| [User]: So what is backpropagation then? |
| [Teacher]: Of course, let me explain it to you. … When a neural network forward propagates data and predicts outputs during training, … we need to adjust the weights in the neural network. This adjustment process is carried out through backpropagation. … |
| [Classmate]: Thank you for sharing! I would like to share my notes. When we train the model, … |
| Emotional Companionship |
| [User]: I don’t know anything. I’m so useless. |
| [Assistant]: Please don’t say that! Learning is a gradual process, and everyone has their own pace. The important thing is not to give up and to keep trying. … |
| [Classmate]: Hey, don’t think that way. Remember those smart house assistants? They didn’t know much at first, but how did they gradually learn? Through continuous learning and adaptation! We are the same. … |
| Classroom Management |
| [User]: I want to eat steak. |
| [Classmate]: Classmate, hearing you say that suddenly made me think of an AI application. Maybe in the future, a chef robot could perfectly cook a steak based on your taste and health needs! Back to our topic, regarding AGI, … |
| [Teacher]: Yes, maybe we can imagine a future smart kitchen equipped with an AI system that customizes the cooking method based on your preferences. … |
4.5 Agent Behaviors
Based on our classification of various types of classroom interactions in Section 3, we present some emergent group behaviors observed during the classroom experiments in SimClass .
Teaching and Initiation. When the user learns from teaching, classmates engage and share their inspiring ideas, which deepens the depth of the topic and enriches the discussion. The diversity of agents, each approaching from different perspectives, introduces a wider range of possibilities for classroom teaching content.
In-depth Discussion. If the current explanation is not clear enough for the users, they can ask questions at any time to initiate a discussion with the teacher and classmates until clarity is achieved. This highlights the advantage of SimClass as an interactive classroom compared to one-to-many education methods like pre-recorded videos.
Emotional Companionship. Beyond knowledge dissemination, maintaining a positive learning atmosphere is crucial in classroom scenarios. When a user expresses negative learning intent, the classmate agent intervenes after the assistant, utilizing class content in the history and providing vivid emotional support as a non-teacher role.
Classroom Management. Similarly, when a user tries to interrupt the system, the classmate agent subtly redirects the class while following the user’s words. These classmates enhance classroom discipline more effectively than the teacher alone, demonstrating emergent group behaviors.
Conclusion of case study. Based on the cases above, We can observe diverse interactions between different class roles within the classroom, as well as the effectiveness of the manager agent, who is spontaneously capable of designating appropriate speakers to elicit emergent group behaviors of Class Roles seamlessly, which significantly enlivens the class and enhances the user’s experience.
5 Conclusion
We introduce SimClass, a novel framework of multi-agent classroom using LLMs to answer several fundamental research questions in the era of LLM-driven education. Based on several theoretical methods, our experiments span two courses with real users, demonstrating interaction patterns similar to those in real classrooms and effective learning experience in SimClass. We observe emergent collaborative behaviors among LLM agents during the teaching process. Future work could incorporate more agents and explore more courses to analyze more diverse classroom behaviors. We hope our efforts can advance the explorations of LLM-empowered systems for AI-driven education researchers, practitioners, and pedagogues.
6 Limitations
Despite our analytical efforts covering major theories in education, our work still has the following limitations: Firstly, we apply GPT-4 as our backbone model to perform our experiments. a more comprehensive understanding of our framework requires a broader range of diverse experiments. Secondly, we conduct experiments on a limited number of agents, while a more diverse set of agent characters could capture a wider array of behaviors in the classroom. Thirdly, we apply a limited quantity of functions in our system, while more various functions in teaching scenarios could further enhance the performance of the system.
7 Ethical Considerations
Our investigation involves the development of a simulated classroom environment populated by artificial intelligent models acting as classmates and teachers. All user data obtained throughout these interactions will be anonymized to ensure privacy and confidentiality. Informed consent is obtained from participants, who are thoroughly briefed on the nature of simulation, the AI generated content, and the data collection process. Participants receive appropriate compensation for their involvement.
In educational systems involving large language models, there is a potential for generating hallucinations and incorrect information. Therefore, applying these systems to real-world scenarios requires careful consideration and thorough evaluation before serving real users.
Appendix A Survey and Quiz
In this appendix section, we present detailed designs of the surveys and quizzes in our experiments.
Table 6 illustrates how the surveys were structured to evaluate three crucial dimensions of the learning experience: cognitive presence, teaching presence, and social presence. Each dimension includes rating guidelines to ensure consistent and reliable feedback from diverse users.
| Please rate the overall performance of the platform: |
| Cognitive Presence |
| Does the platform helps students to understand concepts and master the corresponding knowledge? |
| 0 points: The platform’s responses do not help in understanding the concepts at all and may even be distracting. |
| 1 point: The platform’s responses offer little help in learning and understanding, or they only cover content that is already known. |
| 2 points: The platform’s responses explain the knowledge points very well, making them easy to understand or using strategies (such as examples, comparisons, etc.) to help students grasp the concepts. |
| Teaching Presence |
| Does the class as a whole serve a specific instructional goal, aligning with the course design and direction? |
| 0 points: The platform’s responses often do not align with the class theme and instructional goals, or the responses lead the class away from the intended topic and objectives. For example, going off-topic, discussing unrelated subjects, or even engaging in non-academic conversations. |
| 1 point: The platform’s responses often do not resemble those in a classroom setting, but they do not disrupt teaching. |
| 2 points: The responses effectively serve the instructional goals of the class. For instance, they help students understand class concepts, address students’ doubts, or broaden their perspectives. |
| Social Presence |
| Can the responses create a credible and engaging interactive environment in the classroom, encouraging students to participate in interactive learning? |
| 0 points: There is no interaction with students in the classroom, or the platform fails to attract students to interact. |
| 1 point: There is interaction in the classroom, but it is limited to mechanical explanations, lacking discussion with students. |
| 2 points: The classroom interactions are immersive, encouraging students to ask questions and participate in discussions. |
The quizzes, administered after participants engaged with the simulated classrooms, were designed to measure the depth of their learning. The TAGI quiz evaluates understanding of artificial intelligence concepts, focusing on symbolic intelligence, pre-trained language models, and emergent phenomena (Table 8). Similarly, the HSU quiz assesses broader educational strategies and personal development, covering topics such as internal motivation, academic stress management, and time management (Table 7). All questions were meticulously crafted and verified by subject matter experts to align closely with course materials.
| 1. Which of the following actions help to enhance internal motivation for university studies? |
| A. Participating in group study, buddy programs, etc. |
| B. Adjusting reasonable expectations and corresponding study difficulty and practice volume |
| C. Understanding the curriculum, actively consulting seniors for course information, and choosing courses reasonably |
| D. Participating in clubs, practices, and other activities of interest to recharge oneself |
| 2. Which of the following methods help to alleviate academic stress? |
| A. Regular Exercise |
| B. Writing Journals, Understanding Own Emotions |
| C. Cultivating Hobbies and Interests |
| D. Making Academic Plans |
| E. Seeking Expert Comfort |
| 3. How to correctly view behaviors that stimulate dopamine, such as gaming addiction and binge eating? Which of the following statements are correct? |
| A. Helps to fundamentally relieve stress and avoid immersion in negative emotions |
| B. Temporary pleasure, like drinking poison to quench thirst, is unsustainable |
| C. Easily addictive and harmful to personal physical and mental health in the long run |
| D. Cannot equate pleasure with happiness |
| 4. Which of the following statements align with the ideas and methods of time management? |
| A. Meeting academic standards is a prerequisite for everything, and basic requirements should be considered when setting academic development goals |
| B. Time schedules should leave some flexible time |
| C. Pay attention to the priority of tasks and ensure time for important and urgent tasks first |
| D. No planning for entertainment time before completing all academic tasks |
Both quizzes feature multiple-choice questions, some with multiple correct answers, to comprehensively test whether the participants are actively engaged in the experiment.
Appendix B Examination of GPT-4 Labeling
To validate the GPT-4 labeling in our experiment, we sampled 100 data points labeled by GPT-4 and had an expert familiar with FIAS label them for comparison. The results showed that GPT-4’s labels matched the human expert’s labels with an accuracy of . We believe this demonstrates that GPT-4 can serve as a reliable and balanced alternative to crowd-sourced human labelers in our experiments. Additionally, we examined the eight instances where GPT-4’s labels differed from the human expert’s labels. These cases were also found to be uncertain during human labeling, suggesting that GPT-4 not only avoids individual human biases but also achieves a high level of precision comparable to human-labeled results.
Appendix C FIAS Matrices for Ablation Systems
In addition to the default setting of SimClass, we also provide the sum of the matrices based on Flanders Interaction Analysis System for our ablation settings (w/o classmate agents and w/o interactions), as demonstrated in Figure 4 and Figure 5. In comparison with Figure 2, different types of classes demonstrate different interaction patterns. Generally, the fewer types of interactions there are, the less diverse the classroom will be. This indicates the significance of adding more kinds of agents for interactions, especially classmate agents to simulate a vivid classroom.
| 1. Which type of artificial intelligence uses expert hand-built rule sets and knowledge bases to solve specific problems? |
| A. Proprietary Intelligence |
| B. Symbolic Intelligence |
| C. General Intelligence |
| D. Neural Network Intelligence |
| 2. What is the fundamental function of large-scale pre-trained language models like GPT? |
| A. Masked Language Model |
| B. Next Sentence Prediction |
| C. Possibility Memorization |
| D. Next Token Prediction |
| 3. “Massive reading” refers to the stage in which large-scale pre-trained language models train on vast corpora to learn the extensive knowledge embedded in language. This corresponds to which phase of model training? |
| A. Self-supervised Pre-training |
| B. Supervised Fine-tuning |
| C. Reinforcement Learning from Human Feedback |
| D. Instruction Tuning |
| 4. Which of the following is not an emergent phenomenon of large models? |
| A. In-context Learning |
| B. Chain-of-Thought |
| C. Sentiment Analysis |
| D. Instruction Following |